### v1

# Accretion vs. Core-Filament Collision

Abstract
Context. Recent millimeter and submillimeter observations have unveiled elongated and asymmetric structures around protostars. These structures, referred to as streamers, often exhibit coherent velocity gradients, seemingly indicating a directed gas flow towards the protostars. However, their origin and role in star formation remain uncertain.
Aims. A protostellar core Per-emb-2, located in Barnard 1, has a relatively large streamer with au, which is more prominent in emission from carbon chain molecules. We aim to unveil the formation mechanism of this streamer.
Methods. We conducted mapping observations towards Per-emb-2 using the Nobeyama 45-m telescope. We targeted carbon chain molecular lines such as CCS, HC3N, and HC5N at 45 GHz.
Results. Using astrodendro, we identified one protostellar and four starless cores, including three new detections, on the Herschel column density map. The starless and protostellar cores are more or less gravitationally bound. We discovered strong CCS and HC3N emissions extending from the north to the south, appearing to bridge the gap between the protostellar core and the starless core north of it. This bridge spans au with the velocities from 6.5 to 7.0 km s-1. The bridge has the velocity gradient opposite to the streamer. Thus, the streamer is unlikely to be connected to this bridge, suggesting that the streamer does not have an accretion origin.
Conclusions. We propose that a collision between a spherical core and the filament has shaped the density structure in this region, consequently triggering star formation within the head-tail-shaped core. In this core-filament collision (CFC) scenario, the collision appears to have fragmented the filament into two structures. The streamer is a bow structure, while the bridge is a remnant of the shock-compressed filament. Thus, we conclude that the Per-emb-2 streamer does not significantly contribute to the mass accumulation towards the protostar.
Key Words.:
stars: formation – ISM: clouds – ISM: structure – submillimeter: ISM1 Introduction
Stars form from dense cores in molecular clouds. According to the standard scenario of star formation (Shu et al., 1987), a nearly spherical dense core undergoes gravitational contraction, forming a nearly axi-symmetric system consisting of a protostar and a rotationally-supported disk. However, recent observations have uncovered significant non-axisymmetric structures surrounding young protostellar systems, commonly referred to as streamers or spirals (Pineda et al., 2020; Sanhueza et al., 2019; Chen et al., 2021; Valdivia-Mena et al., 2022; Hsieh et al., 2023; Olguin et al., 2023). Hereafter, we will use the term ”streamer” to refer to such structures for the sake of simplicity. These structures often exhibit coherent velocity patterns, interpreted as infall towards the central protostar (Pineda et al., 2020; Sanhueza et al., 2021). Understanding the role of streamers in mass loading towards central stars is crucial for unveiling the complexities of stellar mass accumulation process.
Numerical studies have proposed various mechanisms for streamer formation. For example, in many numerical simulations of cluster formation, the intermediate-density structure connected with dense cores and sink particles, which resemble the streamers, are often observed and these structures partly play a role in mass accretion onto the central protostellar system (e.g., Bonnell & Bate, 2006; Nakamura & Li, 2007; Kuffmeier et al., 2017; Padoan et al., 2020). On the other hand, Tu et al. (2024) demonstrated that streamer-like structures could arise from local turbulent compression within magnetized parent cores, without the need for external accretion of ambient gas. In this case, the protostellar system does not have to have structures connected with the inter-core medium, and streamers can be generated through the accretion process within the natal core. Therefore, the accretion of ambient gas is likely to be a minor process for supplying the gas, and the stellar mass is roughly determined by the parent core mass. Yano et al. (2024) proposed that streamers could form through dense core collisions (DCC, see also Kinoshita & Nakamura, 2022). They estimated the frequencies of DCC in nearby star-forming molecular clouds and found that the typical dense core experiences at least a collision in its lifetime, particularly in clustered environments. In this DCC model, the compressed layer created between two cores evolves into streamers and spirals, creating the rotating envelope around the protostellar system. In this case, the streamer contributes to supply additional mass into the central protostar, but the mass accretion occurs temporarily, depending on the conditions such as the impact parameter and collision speed. However, these mechanisms await further elucidation based on observational evidence.
To shed light on the streamer formation, we conducted observations towards Per-emb-2, a core with a streamer towards the central protostar (Pineda et al., 2020). It is located in the southern part of Barnard 1 in the Perseus molecular cloud at a distance of 300 pc from the Sun (Zucker et al., 2019). Gas within the streamer appears to exhibit a velocity gradient towards the central protostar [from 7.5 km s-1 (far side) to 7.0 km s-1 (protostar)], and the streamer is particularly prominent in emissions from chemically early-phase molecules like CCS and HC3N as observed with the NOEMA interferometer. Since these molecules steeply decrease with time at the high-density environment, Pineda et al. (2020) claim that the gas of the streamer is supplied from outside via accretion. Very recently, Taniguchi et al. (2024) also claimed that the ambient gas may be provided from the northern part of the protostar.
2 Observations and data
We carried out the on-the-fly (OTF) mapping observations towards the protostellar core Per-emb-2 using the Z45 receiver of the NRO 45m telescope (Nakamura et al., 2014) in the 2022–2023 season. Some details of the observations are presented in the Appendix. The mapped area is determined such that the protostar is located at the centre of the box (Fig. 1a).
The target lines are CCS (), HC3N (), and HC5N (). The parameters are summarised in Table 2. The telescope has a full-width at half-maximum (FWHM) beam size of 38″ at 43 GHz. The standard chopper wheel method was used to convert the output signal into the antenna temperatures () and corrected for the atmospheric attenuation. The main beam efficiency of Z45 is at 45 GHz. The main-beam temperature was calculated as . We adopted a spheroidal function as a gridding convolution function to calculate the intensity at each grid point of the final cube data with a spatial grid size of 15″ and a frequency resolution of 3.81 kHz which corresponds to km s-1 at 45 GHz. The final effective angular resolution of the map was 49″, corresponding to 15000 au, or 0.07 pc at the distance of 300 pc (Zucker et al., 2019).
We downloaded the high resolution H2 column density and dust temperature maps obtained by the Herschel Gould Belt Survey (André et al., 2010; Pezzuto et al., 2021). The angular resolution of the maps are 18″.
Number | Peak R.A. | Peak Dec. | ) | type / HGBS | ||||||
(J2000) | (J2000) | pc | cm-2 | cm-3 | K | km s-1 | ||||
1 | 3:32:19.13 | 30:51:44.67 | 0.508 | 0.0148 | 1.28 | 5.4 | 12.7 | 0.37 | 1.422 (0.565) | starless/449 |
2 | 3:32:16.36 | 30:51:08.40 | 0.398 | 0.0077 | 1.25 | 29.6 | 12.6 | 0.41 | 1.014 (0.377) | starless/new |
3 | 3:32:18.50 | 30:49:50.63 | 9.723 | 0.0884 | 1.97 | 0.5 | 12.1 | 0.84 | 1.077 (0.176) | protostellar/447 |
4 | 3:32:26.94 | 30:48:21.45 | 0.200 | 0.0130 | 0.64 | 3.1 | 13.8 | – | – (1.265) | starless/new |
[…]
HGBS 442 | 3:32:12.04 | 30:47:59.0 | 0.105 | 0.0300 | 0.25 | 0.05 | 0.02 | 10.4 | (3.9) | starless |
3 Results and analysis
3.1 Dense core identification and their dynamical states
In Fig. 1a, we present the Herschel column density map of the observed area. We applied astrodendro (Rosolowsky et al., 2008) to this image and identified five cores, four cores fully covered by our observation box (see the Appendix B.1 for more details). The boundaries of the identified cores are outlined in red. Their physical quantities are summarised in Table 1. In the same area, Pezzuto et al. (2021) identified three cores using getsources: a protostellar core (HGBS 447) and two starless cores (HGBSs 442, 449). Our core Nos. 1 and 3 correspond to HGBS 442 and 447, respectively. The cores Nos. 2, 4 and 5 are starless and newly detected. We could not identify HGBS 449, which has the lowest column density. The central protostellar core (No. 3) has a mass of . According to Pineda et al. (2020), the protostellar core has a systemic velocity of km s-1 and has a streamer with a length of au (Pineda et al., 2020).
From the CCS cube data, we derived the core’s CCS velocity dispersion () and then converted it to the intrinsic velocity dispersion () of gas with a mean molecular weight of to assess their dynamical state. See the Appendix B.2 for the details of the analysis. The cores Nos. 1, 2, 3, and 5 have , where the virial parameter was computed under the assumption of centrally-condensed () sphere. The protostellar core has a significantly-small Bonnor-Ebert ratio with 0.18, but is close to unity, where the Bonnor-Ebert ratio is the ratio between the Bonnor-Ebert critical mass of the core and the observed mass. The core No. 4 does not have significant CCS emission, and we could not measure the velocity dispersion, but the Bonnor-Ebert ratio of the core No. 4 ( 1.3) suggests it is close to bound. We also calculate the external pressures exerted on the core surfaces, and found that they play a minor role except for No. 5, where we calculated the core’s external pressure () from the ambient density and ambient gas intrinsic velocity dispersion measured in the branch just below the corresponding core. The branch is an intermediate structure identified by astrodendro. The core No.5 is dynamically compressed by the ambient pressure and may be hard to create from spontaneous gravitational fragmentation for which cores become self-gravitating. The other cores are close to be in dynamical equilibrium.
For the protostellar core, the column density distribution is not symmetric, but exhibiting a head-tail shape. The centre of gravity determined from this distribution exhibits a significant offset from the position of the protostar itself (see Fig. 8). This suggests that the streamer does not point directly towards the gravity centre of the protostellar core. If its curved structure is generated by the angular momentum of the infalling gas, a large ( au) rotating envelope should form around the protostar and connect to the streamers. However, there is currently no observational evidence for such a large structure (e.g., Tobin et al., 2016; Pineda et al., 2020). Therefore, it is unlikely that the streamer is driven by gravitational infall from outside the core.
3.2 CCS/HC3N bridge with a length of au (0.2 pc)
In Fig. 1, we present the velocity integrated maps of CCS, HC3N, and HC5N alongside with the Herschel column density map. In the CCS integrated intensity map, we recognise the two prominent CCS elongated features: an elongated feature, stretching from north to south, bridging the gap between the starless cores Nos. 1&2 and the protostellar core No. 3, while another compact emission is positioned to the west of the protostellar core.
The distribution of HC3N emission appears to resemble that of CCS, with strong concentrations at the southern periphery of the starless cores Nos. 1&2. Another significant peak is evident immediately above the protostar, with a fainter peak located to the west, coinciding with a local maximum in CCS intensity. HC5N, exhibiting the weakest intensity among the three lines, has a peak at a similar position of the second peak of HC3N, or the northern portion of the protostellar core.
Figs. 2 & 7 display the velocity channel maps of CCS, HC5N, and HC3N. The starless cores Nos. 1&2 appear associated with relatively strong compact emission around a velocity of 6.5 km s-1.
The CCS emission within the velocity range of 6.5–6.7 km s-1 exhibits a fragmented filamentary structure, extending along a line from the northeast to the southwest which coincides with the large-scale filament seen in the H2 column density image (see Fig. 8). An elongated structure is discernible in the southeast corner of the CCS map, though the peak is faint in the H2 map (see Fig. 1).
At a velocity of 6.6–7.0 km s-1, the elongated structure observed in CCS gradually connects with the central protostellar core, forming a bridge. A similar bridging structure is also observed in the corresponding velocity channel for HC5N. Overall, these emissions demonstrate a weak velocity gradient spanning from 6.5 km s-1 to 7.0 km s-1 towards the protostar.
3.3 Mismatch in velocity structure between bridge and streamer
Based on the observations detailed in the previous section, it becomes evident that the protostellar core No. 3 and the northern starless cores Nos.1&2 exhibit a connection characterized by a bridge structure. In Fig. 3, we present the CCS position-velocity plot tracing the 0.2 pc bridge structure.
The protostellar core No. 3 exhibits a velocity of 7 km s-1, indicating a velocity disparity between the two cores, bridged by an intermediate velocity region. The velocity gradient along the bridge is estimated to be 0.5 km s-1/0.2 pc 2.5 km s-1 pc-1. The existence of the intermediate-velocity gas bridging the two structures with different velocities suggests a dynamical interaction between the cores, a similar feature to the larger-scale cloud-cloud collisions (e.g., Nakamura et al., 2012; Nguyen-Lu’o’ng et al., 2013; Fukui et al., 2021; Kinoshita et al., 2021). The enhanced CCS emission observed could potentially stem from the colliding region, where CCS formation may be accelerated due to increased density. Both CCS and HC3N exhibit critical densities around cm-3, implying that while the shocked layer may possess larger density, the column density might not be sufficiently high to be recognised in the 2D Herschel image.
The streamer seemingly follows the cm-2 contour of the protostellar core and is spatially overlapped with the distribution of the CCS emission (see Fig. 2), but the velocities are very different. The streamer has much steeper velocity gradient in the opposite direction to that of the bridge (see the right panel of Fig. 3). Therefore, we conclude that the bridge is not a structure connected to the streamer.
4 Discussion
4.1 Accreting streamer or not?
Taniguchi et al. (2024) claimed that the streamer connects to the northern region through the bridge based on the integrated intensity map. They suggested that the streamer is accreting flow. However, as mentioned above, the gas in the bridge is likely to be different component from the streamer since they have different velocities: the bridge has the velocities from 6.5 (north) to 7.0 km s-1 (south) with the velocity gradient of 2.5 km s-1 pc-1 (hereafter, we define this direction is positive), whereas the streamer has the velocities from 7.5 (north) to 7.0 km s-1 (south) with the velocity gradient of –0.5 km s-1/0.05 pc = –10 km s-1 pc-1. There is no evidence showing that the streamer connects to the bridge in the velocity distribution.
In addition, the centre of gravity of the protostellar core has a significant offset from the protostar of 12″( au). If the streamer originates from the inflow from the ambient gas driven by the gravity, it is likely to point towards the core’s gravity centre (see Fig. 8). Or, if the offset is due to rotation, this object must have a large rotating envelope with a size of au. Such a structure is not detected towards the protostar (Tobin et al., 2016). This fact also implies that the streamer is not accretion origin, contrasting with the interpretation of Pineda et al. (2020). The northern cores (1&2) are close to gravitationally bound and therefore there is a possibility for them to evolve into protostars. Their merging might trigger star formation. The protostellar core also contains appreciable mass, a part of which may eventually accrete onto the central protostar before the ambient gas falls in.
Yano et al. (2024) proposed the DCC scenario to reproduce the density structure in Per-emb-2 in numerical simulations. According to this model, when two cores collide with each other, the shock compressed layer is formed in between, connected to the protostar. The gas is inflowing towards the protostar. However, if this is the case, both the bridge and streamer should have the velocity gradients with the same directions (see Kinoshita & Nakamura, 2022). The velocity/density structure appears more complicated than what the DCC model predicts. Therefore, the formation mechanism of the streamer is neither ambient gas inflow nor two round core collision.
4.2 A core-filament collision (CFC) model
From the evidence of the existence of the large scale filament, the positive gradient towards the streamer of the protostar. We propose a core-filament collision (CFC) model to explain for the configuration of Per-emb-2, Fig. 4 depicts the schematic representation of this CFC model.
We propose that (1) a spherical core, with a relative speed of 1–1.5 km s-1, approached from the west side of the filament and collided with it. (2) The collision initiated star formation on the far side of the eastern extremity of the core, where the initial shock compression occurred, leading to the formation of a protostellar system. The arc-like streamer may be the bow structure formed by the CFC, and therefore the velocity gradient of the streamer is not due to the gravitational accretion. If its dynamics is not significantly affected by the stellar gravity, the flow is pointed towards outside along it (outflow), instead of infall. The bow typically has an almost linear velocity structure along it because of its curvature. Such a feature is consistent with the observation. When the stellar gravity is sufficiently strong to attract the gas towards the star, the flow may be pointing towards the protostar. It would be difficult to judge which direction (outflow or inflow) is reasonable from the current observational data. Future numerical experiments would be needed to verify this scenario more quantitatively.
The head-tail morphology observed in the No. 3 core agrees well with the anticipated appearance resulting from this collision. The CCS arc (streamer) structure (Pineda et al., 2020) appears to well trace the edge of the protostellar core ( cm-2, see Fig. 2).
Furthermore, the collision led to the division of the filament into two distinct structures (Nos. 1&2 and 5). The shock-compressed layers in the northern and southern regions attained densities of cm-3, resulting in relatively strong CCS/HC3N emissions. The northern shock-compressed layer became interconnected with the central protostellar system. There is a possibility that it could evolve into a accreting streamer in the future. However, the total accretion mass via the bridge (future streamer) may be small. The protostellar core envelope still contains enough material (). Therefore, the accretion along the streamers, if it exits, plays only a minor role in the mass accumulation towards the central protostar.
The total velocity difference is about 1 (= 7.5–6.5) km s-1 and the filament width is about 0.1 pc, therefore it takes 0.1 pc / 1 km s-1 yr for the core to cross the filament. This is comparable to the typical lifetime of Class 0 protostar and consistent with the triggered star formation scenario proposed by CFC. Furthermore, if the bridge is the remnant of the shock compressed layer, its density increased from cm-3 to cm-3 by the isothermal shock with a Mach number of 1.5 km s-1 / sound speed . This condition in the shocked area, or streamer, is quite similar to the condition assumed by Suzuki et al. (1992) who started the chemical evolution calculation assuming a steady state with a uniform isothermal gas of cm-3 and K. According to Fig. 13 of (Suzuki et al., 1992), the CCS/HC3N becomes most abundant in yr in such a density range, and declines steeply after 0.5 Myr (see also Leung et al., 1984). These considerations on the timescales well match the CFC scenario. In our CFC or DCC scenarios, shocks are likely to somewhat increase gas temperature although the molecular gas is highly radiative and therefore the gas temperature is kept low even after shock compression. In such a situation, the streamers formed by the collision model may have emission of SO or CH3OH whose sublimation temperatures are about 50 and 80 K, respectively. These lines could exhibit broader widths as a result of shock interactions. The above consideration should be validated through future chemical evolution calculations using e.g., the Paris-Durham shock code (Godard et al., 2019).
Acknowledgements.
This work was financially supported by JSPS KAKENHI Grant Numbers JP23H01218 (F.N.). Part of this work was supported by the NAOJ Visiting Research Grant (Q.N.L.). We thank the NRO staff for both operating the 45 m and helping us with the data reduction. We thank the anonymous referee for valuable comments that helped improve the paper.References
- André et al. (2010) André, P., Men’shchikov, A., Bontemps, S., et al. 2010, A&A, 518, L102
- Bertoldi & McKee (1992) Bertoldi, F. & McKee, C. F. 1992, ApJ, 395, 140
- Bonnell & Bate (2006) Bonnell, I. A. & Bate, M. R. 2006, MNRAS, 370, 488
- Chen et al. (2021) Chen, X., Ren, Z.-Y., Li, D.-L., et al. 2021, ApJ, 923, L20
- Fukui et al. (2021) Fukui, Y., Habe, A., Inoue, T., Enokiya, R., & Tachihara, K. 2021, PASJ, 73, S1
- Godard et al. (2019) Godard, B., Pineau des Forêts, G., Lesaffre, P., et al. 2019, A&A, 622, A100
- Hsieh et al. (2023) Hsieh, T. H., Segura-Cox, D. M., Pineda, J. E., et al. 2023, A&A, 669, A137
- Kinoshita & Nakamura (2022) Kinoshita, S. W. & Nakamura, F. 2022, ApJ, 937, 69
- Kinoshita et al. (2021) Kinoshita, S. W., Nakamura, F., Nguyen-Luong, Q., et al. 2021, PASJ, 73, S300
- Kuffmeier et al. (2017) Kuffmeier, M., Haugbølle, T., & Nordlund, Å. 2017, ApJ, 846, 7
- Leung et al. (1984) Leung, C. M., Herbst, E., & Huebner, W. F. 1984, ApJS, 56, 231
- Maruta et al. (2010) Maruta, H., Nakamura, F., Nishi, R., Ikeda, N., & Kitamura, Y. 2010, ApJ, 714, 680
- Nakamura et al. (2024) Nakamura, F., Chiong, C.-C., Taniguchi, K., et al. 2024, PASJ, 76, 563
- Nakamura & Li (2007) Nakamura, F. & Li, Z.-Y. 2007, ApJ, 662, 395
- Nakamura et al. (2012) Nakamura, F., Miura, T., Kitamura, Y., et al. 2012, ApJ, 746, 25
- Nakamura et al. (2014) Nakamura, F., Sugitani, K., Tanaka, T., et al. 2014, ApJ, 791, L23
- Nguyen-Lu’o’ng et al. (2013) Nguyen-Lu’o’ng, Q., Motte, F., Carlhoff, P., et al. 2013, ApJ, 775, 88
- Olguin et al. (2023) Olguin, F. A., Sanhueza, P., Chen, H.-R. V., et al. 2023, ApJ, 959, L31
- Padoan et al. (2020) Padoan, P., Pan, L., Juvela, M., Haugbølle, T., & Nordlund, Å. 2020, ApJ, 900, 82
- Pezzuto et al. (2021) Pezzuto, S., Benedettini, M., Di Francesco, J., et al. 2021, A&A, 645, A55
- Pineda et al. (2020) Pineda, J. E., Segura-Cox, D., Caselli, P., et al. 2020, Nature Astronomy, 4, 1158
- Rosolowsky et al. (2008) Rosolowsky, E. W., Pineda, J. E., Kauffmann, J., & Goodman, A. A. 2008, ApJ, 679, 1338
- Sanhueza et al. (2019) Sanhueza, P., Contreras, Y., Wu, B., et al. 2019, ApJ, 886, 102
- Sanhueza et al. (2021) Sanhueza, P., Girart, J. M., Padovani, M., et al. 2021, ApJ, 915, L10
- Shu et al. (1987) Shu, F. H., Adams, F. C., & Lizano, S. 1987, ARA&A, 25, 23
- Suzuki et al. (1992) Suzuki, H., Yamamoto, S., Ohishi, M., et al. 1992, ApJ, 392, 551
- Taniguchi et al. (2024) Taniguchi, K., Pineda, J. E., Caselli, P., et al. 2024, ApJ, 965, 162
- Tobin et al. (2016) Tobin, J. J., Looney, L. W., Li, Z.-Y., et al. 2016, ApJ, 818, 73
- Tu et al. (2024) Tu, Y., Li, Z.-Y., Lam, K. H., Tomida, K., & Hsu, C.-Y. 2024, MNRAS, 527, 10131
- Valdivia-Mena et al. (2022) Valdivia-Mena, M. T., Pineda, J. E., Segura-Cox, D. M., et al. 2022, A&A, 667, A12
- Yano et al. (2024) Yano, Y., Nakamura, F., & Kinoshita, S. W. 2024, ApJ, 964, 119
- Zucker et al. (2019) Zucker, C., Speagle, J. S., Schlafly, E. F., et al. 2019, ApJ, 879, 125
Appendix A Observations and target lines
The OTF observations were carried out in October 2022 (2 nights) and in March 2023 (1 daytime). Z45 is dual linear polarization receiver and we summed up the horizontal and vertical polarization components to improve the signal-to-noise ratios. The scan interval of the OTF observations was set to 8″, about a fifth of the beam size. The pointing observations were made every one hour with SiO maser lines of NML-Tau (KL Tau). The pointing errors were within 3″– 5″. We used SAM45 digital spectrometer as a backend. SAM45 is a highly flexible FX-type digital spectrometer with a finest frequency resolution of 3.81 kHz. The bandwidth of SAM45 with 3.81 kHz was 15.625 MHz. It took about 30 min to obtain a single map. We summed 6 maps to obtain a final map. The total observation time was about 6 hours including the overhead of pointing observations. The on-source time was 3 hours. During the observations, the typical system noise temperatures were about 150–200 K.
The rest frequencies and transitions of the target lines are summarised in Table 2. See Nakamura et al. (2024) for the rest frequency of CCS.
Species | Transition | Frequency |
- | - | (GHz) |
CCS | 45.379033 | |
HC3N | 45.490316 | |
HC5N | 45.264720 |
Appendix B Core identification and physical quantities
B.1 Core identification
We applied astrodendro (Rosolowsky et al. 2008) to the high-resolution Herschel column density map with an angular resolution of 18″. The astrodendro has three parameters for the hierarchical structure analysis: min_value which is the threshold value, min_delta which is the minimum step for structure identification, and min_npix which is the minimum pixel number. We adopt min_value , min_delta , and min_npix = Beam size/pixel size, cm-2) is the rms noise level. The astrodendro identifies three structures: a leaf, branch, and trunk. The minimum structure of “a leaf” is defined as a dense core. To estimate the ambient gas pressure around the cores, we adopted the physical quantities averaged in a branch immediately below the corresponding leaf. We also derive the core mass by subtracting the background component which is defined as the trunk.
B.2 Dynamical states of the cores
We assess the dynamical states of the cores applying the virial theorem. The virial equation for a spherical core is given by
where , , , and are the moment of inertia, internal kinetic energy, gravitational energy, and surface pressure term, respectively, and the magnetic field terms are neglected (Maruta et al. 2010). These individual terms are given as
where , , , and are the gravitational constant, core mass, core radius, and the external pressure, respectively. is a dimensionless parameter of order unity which measures the effects of a non-uniform or nonspherical mass distribution (Bertoldi & McKee 1992). For a uniform sphere and a centrally condensed sphere with , = 1 and 5/3, respectively. Here, we adopt since the cores tend to be more or less centrally-condensed. is the 1D intrinsic velocity dispersion of the molecule of mean mass,
and is the gravitational constant, is the Boltzman constant, is the velocity dispersion measured by CCS, (=2.33) is the mean molecular weight, is the mass of a hydrogen atom, and is the mass of a CCS molecule. The CCS velocity dispersions towards the cores are computed by the Gaussian fitting of the averaged spectra (see Fig. 5). It is worth noting that CCS emission may preferentially trace the outer envelopes of cores, given its critical density of cm-3. In contrast, the core densities calculated from the Herschel image range from to cm-3 as shown in Table 1. Despite this, in the core analysis presented here, we adopt the CCS line widths as representative of the cores’ properties.
The external pressure is calculated as
where and are the density and 1D intrinsic velocity dispersion of ambient gas, respectively, and they are derived from the values averaged in the branch immediately below the corresponding leaf. The ambient density was computed under the assumption that the branch is spherical symmetry.
The core virial parameter is calculated as
Table 3 summarises the velocity dispersions of the cores and their energy ratios. Fig. 6 shows the energy ratios vs. , where W, U, and S are the gravity term, internal kinetic and thermal energy term, and the surface term, respectively. In this plot, the cores below the black dashed line are satisfied with the condition , for which the cores collapse. The core’s self-gravity is important for the area above the dotted line (), whereas the surface pressure is dominant in the area below the dotted line. The cores 1, 2, and 3 are located in the gravity-dominant area. They are close to the dynamical equilibrium and self-gravitating. For the core 5, the external pressure plays a crucial role in the core’s dynamical stability, and thus the core is dynamically unstable due to the external pressure. Such a core is difficult to create from the gravitational fragmentation of the filament. The formation of unstable non-self-gravitating core (core No. 5) appears to agree with the collision scenario.
Appendix C Channel maps
Fig. 7 shows the channel maps of HC3N ( and HC5N (). The HC3N has a stronger emission but having several overlapped hyperfine components. However, adopting the rest frequency of as a centre frequency, we simply integrated all the emission to construct the channel maps. The bridge is prominent for HC3N (, and structure similar to CCS was detected. For HC5N (, the strong emission is seen just at the southern edge of cores No. 1 and 2.
Number | ||||||||
(km s-1) | (km s-1) | (km s-1) | (cm-3) | (km s-1) | ||||
1 | 0.37 | 0.41 | 0.26 | 0.28 | -0.703 | -0.111 | 1.422 | |
2 | 0.41 | 0.41 | 0.27 | 0.27 | -0.986 | -0.019 | 1.014 | |
3 | 0.84 | 0.61 | 0.41 | 0.33 | -0.928 | -0.734 | 1.077 | |
5 | 0.45 | 0.61 | 0.28 | 0.34 | -0.339 | -3.025 | 2.954 |
Appendix D Centre of gravity of the protostellar core w.r.t the large filament
The circle in Fig. 8 indicates the position of the protostar. The black square indicates the column density weighted position of the protostellar core. If the gas were distributed only on the plane of sky, the position would coincide with the centre of gravity. In other words, the protostar was not formed at the core centre but near the core envelope. This implies that the core formation did not occur spontaneously.
If the curved shape of the streamer identified by Pineda et al. (2020) is due to the angular momentum of infalling gas, the rotational motion must be significant on the scale of au, leading to the formation of a large ( au) rotating envelope or disk. However, such a large circumstellar structure was not detected in the observations by Tobin et al. (2016). Thus, at this scale, the rotational motion would still be minor and unlikely to create the elongated, curved structure seen in Fig. 2. For example, Yano et al. (2024)’s core collision simulation showed that the streamer formed by the infalling gas from the second core is smoothly connected to a central, rotating disk-like envelope. Such a structure is a plausible outcome of the infalling gas scenario. Since the rotating disk is small for this protostar, the curved structure, if formed by the infalling gas, must almost point directly to the protostar.